# Univariate-Guided Sparse Regression

### v8

@@ 1 Introduction
High-dimensional regression and classification problems are ubiquitous across fields such as genomics, finance, and social sciences. In these settings, the Lasso (Least Absolute Shrinkage and Selection Operator) has emerged as a widely-used methodology due to its ability to simultaneously perform variable selection and regularization, yielding sparse and interpretable models. Despite its widespread use, Lasso has certain limitations, including sensitivity to correlated predictors and inclusion of spurious features in the final model. There have been many related proposals including MC+ (?), the elastic net (?), the adaptive Lasso (?), and SparseNet (?).
@@
Here is an example, taken from unpublished work with collaborators of the third author. (Since the work is unpublished, we do not give background details). The dataset is from cancer proteomics, with 559 biomarkers measured on 81 patients, 20 healthy and 61 with cancer. We divided the data into 70% training and 30% test, and applied both the lasso and uniLasso to the training set. Both methods performed well, giving cross-validation error of about 5%, and 4 and 3 errors respectively out of 30 test samples. Table 1 shows the results of lasso and uniLasso applied to this data, using cross-validation to choose the model. In the top table there are 2 sign changes from univariate to multivariate: #3 goes from mildly protective to a negative indicator, while # 7 goes in the opposite direction.
By design, uniLasso in the bottom table has no sign changes. Further, its univariate coefficients are much larger (in absolute value) than those of the lasso. This makes it a more credible and stable model. We investigate stability in Section 10.1.
The remainder of this paper is organized as follows. Section 2 provides a detailed description of the UniLasso methodology. We examine the relationship between UniLasso and the Adaptive Lasso in Section 3. We compare UniLasso to least squares and Lasso in the well-known diabetes data set in Section 4: here we show the utility of unregularized uniLasso as an alternative to least squares. In Section 5 we look at the orthonormal design case, where we can derive explicit an expression for the UniLasso solution. Section 6 presents some theoretical results on support recovery and mean square error consistency. Section 7 presents simulation studies comparing UniLasso to Lasso and other benchmark methods, while Section 8.1 examines uniReg, the unregularized ($\lambda=0$) case. Section 9 applies UniLasso to real-world datasets, and in Section 10 we present an “adversarial attack” view, and multiclass and GWAS examples. Section 11 explores the setting where additional data is available, not full data but just the univariate coefficients. In Section 12 we study whether cross-validation works properly in the UniLasso setting. We discuss UniLasso for GLMs and the Cox Survival Model, as well as computation in Sections 13 and 14. Lastly, Section 15 concludes with a discussion of future directions.
@@
Biomarker Univariate LS coefficient Lasso coefficient Sign change?
1 -12.709 -1.845
2 25.269 37.492
3 -0.34 6.245 X
4 -24.74 -11.262
5 -17.966 -11.908
6 -49.24 -18.919
7 1.81 -6.265 X
8 22.909 13.995
9 -37.646 -5.415
10 13.024 0.79
11 -10.141 -2.325
12 -22.383 -12.745
@@
Biomarker univariate multivariate
A 58.924 4.982
B 25.269 6.286
C -49.24 -13.667
D 73.844 38.448
E 69.152 23.034
F -54.94 -9.261
G 77.688 9.458
Table 1: Proteomics study. Top table: Univariate least squares coefficients and lasso coefficients, for model chosen by Lasso. Bottom table: same for model chosen by uniLasso. All biomarkers are different in the two tables, with the exception of #2 and #6 in the first table corresponding to B and C in the second.
@@ 2.1 Our proposal
Our procedure has 3 simple steps, which we motivate here. For interpretability and prediction accuracy, we preprocess the features in Step 1, multiplying them by a robust version of the univariate least-squares coefficient estimates. In Step 2 we fit a non-negative lasso. Together these ensure that
@@
    For $j=1,2,\ldots,p$, fit the $p$ separate univariate least-squares models $\hat{\eta}_{j}(x_{j})=\hat{\beta}_{0j}+\hat{\beta}_{j}x_{j}$ to the response $y$. For $j=1,2,\ldots,p$ and each $i\in 1,2,\ldots,n,$ compute the leave-one-out (LOO) fitted value $\hat{\eta}_{j}^{-i}=\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij}$, resulting in a new $n\times p$ feature matrix $F=\{\hat{\eta}_{j}^{-i}\}$.
@@
    Fit the Lasso — with an intercept, no standardization, and non-negativity constraints — to target $y$ and these LOO fits as features, giving coefficients $\hat{\theta}=(\hat{\theta}_{0},\hat{\theta}_{1},\cdots,\hat{\theta}_{p})^{\top}$:
    $\mathop{\rm minimize}_{\theta}\;\left\{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\theta_{0}-\sum_{j=1}^{p}\theta_{j}\hat{\eta}^{-i}_{j})^{2}+\lambda\sum_{j=1}^{p}|\theta_{j}|\right\}\quad\mbox{ subject to $\theta_{j}\geq 0,\;j=1,\ldots,p$}.$ (1)
    and return the composite model
    $\hat{\eta}(x)=\hat{\theta}_{0}+\sum_{j=1}^{p}\hat{\theta}_{j}\hat{\eta}_{j}(x_{j}).$ (2)
    Since each of the constituent models $\hat{\eta}_{j}(x_{j})$ are linear, this final model can be written
    $\hat{\eta}(x)=\hat{\gamma}_{0}+\sum_{j=1}^{p}\hat{\gamma}_{j}x_{j},$ (3)
    with $\hat{\gamma}_{j}=\hat{\theta}_{j}\hat{\beta}_{j}$, and $\hat{\gamma}_{0}=\hat{\theta}_{0}+\sum_{\ell=1}^{p}\hat{\theta}_{\ell}\hat{\beta}_{0\ell}$.
This procedure is computationally convenient: Step 1 uses an efficient LOO formula, and then in Step 2, we can apply any efficient $\ell_{1}$ solver. Here we used the R language package glmnet. Specifically, we use the function cv.glmnet to estimate the Lasso path parameter, and have all of the functionality for glmnet at our disposal. We provide a function cv.UniLasso that implements this approach.
Note that we do not standardize the features before applying the non-negative Lasso in Step 2. From our knowledge of multiple linear regression, the first constraint — agreement between univariate and multivariate signs — may seem like an unreasonable restriction. However our belief is that in high-dimensions there are likely to be a multitude of different models that have about the same MSE as the “optimal” one chosen by the Lasso. Hence it can make sense to choose one that is interpretable and sparser than that of lasso.
@@
These conjectures are borne out by our numerical studies. In simulations with varying problems sizes and SNR, and a number of real datasets (see Section 7), in almost every case UniLasso did no worse than the Lasso in terms of out-of-sample MSE, and delivered a substantially sparser model. These real datasets were not “cherry picked” but were the first group of datasets that we tested. We note that ? studies sign-constrained least squares estimation for high-dimensional regression, which relates to our non-negativity constraint in our second step.
Remark A. In equation (1), the first term has a multiplier of $1/n$. In other descriptions of the Lasso, the factor may be set to $1/2$; the glmnet package uses $1/2n$. The different choices lead to equivalent problems and just differ in the scaling of the path parameter $\lambda$. In the theory of Section 6 we assume that the $1/n$ multiplier is used.
@@
Remark B. The UniLasso procedure applies seamlessly to other GLMs as well the Cox model. Indeed, all of the models covered by the glmnet package are at our disposal. All we need are the separate fitted linear models and their LOO fit vectors in step 1, and then glmnet can be directly applied. The only challenge is find an (approximate) LOO formula for the given family. We give details of these computations in the Section 14.
@@
Remark C. UniLasso can be thought of as a version of Stacked regression, or Stacking (??). Stacking is a two-stage procedure for combining the predictions from a set of models (“learners”), in order to get improved predictions. It works as follows. A set of base models is trained on the training data; these models can be of the same type (e.g. gradient boosting) or different types (e.g., linear regression, decision trees, etc.). Each model generates predictions, capturing specific aspects of the relationship between the predictors and the target variable. A meta-model is then trained to combine the predictions of the base models into a final prediction. This meta-model learns how to weigh and integrate the outputs of the base models to minimize the overall prediction error. UniLasso is a special case of stacking where the individual learners are simple univariate regressions.
@@
Remark D. Why do we use the LOO estimates $\hat{\eta}_{j}^{-i}$ as features in step 2, instead of the least squares estimates $\hat{\eta}_{j}^{i}=\hat{\eta}_{j}(x_{ij})$? In traditional stacking this is essential because the individual learners can be of very different complexity. In UniLasso, one would think that the learners (univariate regression) all have the same complexity so that the LOO versions are not needed. Indeed, the theory in Section 6 is based on the LOO estimates but holds equally well for the usual (non-LOO) estimates. However in practice we have found that the LOO estimates often lead to better performance and hence we use them. With the use of the usual (non-LOO) univariate estimates in uniLasso, the resulting estimator is closely related to the adaptive lasso and we explore this connection in Section 3.
@@
Remark E. The non-negative garotte (?) is another closely related method. It minimizes $\sum_{i}(y_{i}-\sum_{j}c_{j}\hat{\beta}_{j}x_{ij})^{2}$ subject to $c_{j}\geq 0$ for all $j$, and $\sum c_{j}\leq s$, where $\hat{\beta}_{j}$ are the usual (multivariable) LS estimates. This is different from our proposal in an important way— our use of univariate LOO estimates in the first step. The univariate coefficients lead to materially different solutions and also allow application to the $p>n$ scenario.
@@ 2.2 UniLasso with no-regularization
We call this “UniReg’ for univariate-guided regression and it represents an interesting alternative to the usual least squares estimates. For example, the non-negative constraint still can provide sparsity even though there is no $\ell_{1}$ penalty. And it gives a unique solution even if $p>n$. The standard error and distribution of the estimated coefficients can be estimated via the bootstrap. We compare UniReg to least squares in the diabetes example of Section 4, and study it in more detail in Section 8.1.
2.3 Two examples
We generated data with 300 observations and 1000 features, the feature covariance having an $AR(1)$ form with $\rho=0.8$. The coefficient vector had 100 non-zero values, uniformly distributed between 0.5 and 2.0, positioned at every other feature. Finally the error variance was 225 yielding an SNR of about 1.5.
The top panels of Figure 1 show the CV and test error curves for the Lasso and UniLasso. We can think of this as a “home court” setting for UniLasso where the pairwise correlations between features is positive, as are the true coefficients. This connection is explained in the theorems of Section 6.
Figure 1: Results for homecourt example. Top panel; CV and Test set prediction error. The dotted vertical lines shows the values of $\lambda$ chosen by CV for each model. Bottom panels: true positives (green) and false positives (red). Horizontal dotted lines: mean value of true nonzero coefficients for the chosen support set (all of which are positive).
We see that UniLasso has test error a little below that of Lasso, with a smaller active set. The bottom panels shows the true positives (green) and false positives (red). UniLasso yields more true positives and fewer false positives.
Table 2 shows the result of 100 simulations from this setup. We see that same pattern emerges. Note that the table shows MSE, rather than PSE as in Figure 1; the two quantities differ in expectation by $\sigma^{2}=225.$ Notice that UniLasso has much higher TPR and much lower FPR than Lasso.
@@ 2.3 Two examples
mean 294.17 282.27 72.00 46.00
se 1.46 1.33 1.78 0.82
TPR-Lasso FPR-Lasso TPR-UniLasso FPR-UniLasso
mean 0.33 0.67 0.49 0.51
se 0.01 0.01 0.01 0.01
Table 2: Test MSE, support, TPR (True Positive Rate) and FPR (False Positive Rate) for 100 simulations from the setting of Figure 1
For the one realization from the same setup but taking $n=300,p=100$, Figure 2 shows the number of nonzero coefficients for lasso and uniLasso as a function of $\lambda$. The right end of the plot represents least squares (for the lasso) and uniReg (for uniLasso). Least squares of course uses all $p=100$ features for $\lambda=0$, but remarkably uniReg uses fewer than 20. It seems that the LOO feature scaling in step 1 and non-negative constraints in step 2 provide implicit regularization.
Figure 2: Lasso and uniLasso: size of support set as $\lambda\rightarrow 0$ for low-SNR setting with $n=300,p=100$.
The next example shows a case where UniLasso performs poorly. Here $n=300,\;p=1000$, and the pairwise correlations between the features and the signs of the 100 true coefficients are both about half positive and negative. The SNR is 4.5. In Section 6 we see that UniLasso does well when the population coefficients for positively correlated features have the same sign, an assumption clearly violated in this example. Fortunately, in any given example, we can assess the relative merits of the Lasso and UniLasso via cross-validation.
Figure 3: Results for simulated high SNR example, where the true number of nonzero coefficients is 100 and $p=1000$. Top panel: CV and test set prediction error. The dotted vertical lines shows the values of $\lambda$, chosen by CV for each model. Bottom panels: true positives (green) and false positives (red).
@@ 3 Relationship of UniLasso to the Adaptive Lasso
The adaptive Lasso (?) is defined by
$\hat{\beta}^{*}={\rm argmin}=\frac{1}{2}\sum_{i=1}^{n}(y_{i}-\beta_{0}-\sum_{j=1}^{p}x_{ij}\beta_{j})^{2}+\lambda\sum_{j=1}^{p}\hat{w}_{j}|\beta_{j}|$ (5)
where $\hat{w}_{j}=1/|\hat{\beta}_{j}|^{\nu}$ and $\hat{\beta}$ is any root-$n$ consistent estimate of $\beta$, for example the least squares estimates.
The original paper assumed $n>p$; with $p>n$, one cannot use the least-squares estimates. One possibility is to use an initial estimate (from say Ridge or Lasso), solve the adaptive Lasso, and iterate. Another option is to use the univariate least-squares estimates: ? consider a combination of these to achieve some theoretical guarantees.
Let $\hat{\beta}_{j,u}$ be the univariate least squares coefficients. Consider the adaptive Lasso with weights $1/|\hat{\beta}_{j,u}|$. Then it easy to show that this procedure is equivalent to uniLasso with $\hat{\eta}_{j}^{i}$ replacing their LOO versions $\hat{\eta}_{j}^{-i}$, and removing the non-negativity constraint in the second stage.
@@
$\min_{\theta}\sum_{i=1}^{n}(y_{i}-\theta_{0}-\sum_{j=1}^{p}\eta_{j}^{i}\theta_{j})^{2}+\sum_{j=1}^{p}|\theta_{j}|\mbox{ s.t. $\theta_{j}\geq 0\;\forall j$}$ (6)
$\min_{\gamma}\sum_{i=1}^{n}(y_{i}-\gamma_{0}-\sum_{j=1}^{p}x_{ij}\gamma_{j})^{2}+\sum_{j=1}^{p}\frac{|\gamma_{j}|}{|\hat{\beta}_{j}|}\mbox{ s.t. sign($\gamma_{j}$)=sign($\hat{\beta}_{j}$) $\forall j$}$ (7)
@@
The proof is in the Appendix. However our experiments with this version of adaptive lasso produced similar results to the one that does not enforce sign constraints,
Here is an example. Figure 4 shows the results from one realization of the “low-SNR” setting of Section 7. We see that both versions of the adaptive lasso overfit, and cross-validation fails to detect this. On the other hand, uniLasso chooses a null model and achieves a much lower MSE. The use of the usual (non-LOO) univariate coefficients in the adaptive lasso causes the problem: it allows the procedure to overfit and cross-validation fails. The reason is that for computational efficiency, cross-validation is not wrapped around both steps (estimation of univariate coefficients and the lasso fit) but is only applied to the lasso fit.
By using LOO fits, uniLasso avoids this problem, because step 1 is internally validated. We note that this poor performance of adaptive lasso does not always occur in this setting, but occurs often enough to hurt its average performance, especially in support size.
Figure 4: CV curves from a low-SNR simulation with $n=300,p=1000$. The test MSE for each chosen model is given above each panel.

### v9

@@ 1 Introduction
High-dimensional regression and classification problems are ubiquitous across fields such as genomics, finance, and social sciences. In these settings, the lasso (Least Absolute Shrinkage and Selection Operator) has emerged as a widely-used methodology due to its ability to simultaneously perform variable selection and regularization, yielding sparse and interpretable models. Despite its widespread use, lasso has certain limitations, including sensitivity to correlated predictors and inclusion of spurious features in the final model. There have been many related proposals including MC+ (Zhang, 2010), the elastic net (Zou & Hastie, 2005), the adaptive lasso (Zou, 2006), and SparseNet (Mazumder et al., 2011).
@@
Here is an example, taken from unpublished work with collaborators of the third author. (Since the work is unpublished, we do not give background details). The dataset is from cancer proteomics, with 559 biomarkers measured on 81 patients, 20 healthy and 61 with cancer. We divided the data into 70% training and 30% test, and applied both the lasso and uniLasso to the training set. Both methods performed well, giving cross-validation error of about 5%, and 4 and 3 errors respectively out of 30 test samples. Table 1 shows the results of lasso and uniLasso applied to these data, using cross-validation to choose the model. In the top table there are two sign changes from univariate to multivariate: #12 goes from mildly protective to a negative indicator, while #11 goes in the opposite direction.
By design, uniLasso in the bottom table has no sign changes. Further, its univariate coefficients are much larger (in absolute value) than those of the lasso. This makes it a more credible and stable model.
The remainder of this paper is organized as follows. Section 2 provides a detailed description of the uniLasso methodology. We examine the relationship between uniLasso and the adaptive lasso in Section 3, and in Section 4 we study the underlying reason for the strong sparsity produced by uniLasso. We compare uniLasso to lasso on a car price data set in Section 5: here we show the utility of unregularized uniLasso as an alternative to least squares. In Section 6 we look at the orthonormal design case, where we can derive an explicit expression for the uniLasso solution. Section 7 presents some theoretical results on support recovery and mean square error consistency. Section 8 presents simulation studies comparing uniLasso to lasso and other benchmark methods, while Section 9 examines uniReg, the unregularized ($\lambda=0$) case. Section 10 applies uniLasso to real-world datasets, and in Section 11 we adapt uniLasso to multiclass classification problems. Section 12 explores the setting where additional data is available: not the complete data, but just the univariate scores. In Section 14 we study whether cross-validation works properly in the uniLasso setting. We discuss uniLasso for GLMs and the Cox Survival Model, as well as computation in Sections 15 and 16. Lastly, Section 17 concludes with a discussion of future directions.
@@
Biomarker Univariate LS Coefficient Lasso Coefficient
1 -49.24 -18.92
2 -37.65 -5.42
3 25.27 37.49
4 -24.74 -11.26
5 22.91 14
6 -22.38 -12.75
7 -17.97 -11.91
8 13.02 0.79
9 -12.71 -1.85
10 -10.14 -2.32
11 1.81 -6.27
12 -0.34 6.25
@@
A 93.13 13.82
B -84.26 -8.17
C 77.69 15.86
D 73.84 27.66
E 69.15 16.21
F 58.92 1.38
G -53.46 -10.5
H -49.24 -20.76
I 25.27 10.48
Table 1: Proteomics study. Top table: Univariate least squares coefficients and lasso coefficients, for model chosen by lasso. The two sign changes from univariate to multivariate are shown in blue. Bottom table: same for model chosen by uniLasso. Biomarkers are ordered by decreasing absolute value. Most biomarkers are different in the two tables, with the exception of 1 and 3 in the first table corresponding to H and I in the second table.
@@ 2.1 Our proposal
Our procedure has three simple steps, which we motivate here. For interpretability and prediction accuracy, we preprocess the features in Step 1, multiplying them by a robust version of their univariate least-squares coefficient estimates. In Step 2 we fit a non-negative lasso tuned by cross-validation, and Step 3 combines the components of Steps 1 and 2 to produce a final model. Together these ensure that:
@@
    For $j=1,2,\ldots,p$ compute the univariate intercepts and slopes $(\hat{\beta}_{0j},\hat{\beta}_{j})$ and their leave-one-out (LOO) counterparts $(\hat{\beta}_{0j}^{-i},\hat{\beta}_{j}^{-i}),\;i=1,\ldots,n.$
@@
    Fit the lasso — with an intercept, no standardization, and non-negativity constraints — to target $y$ and the univariate LOO fits as features
    $\mathop{\rm minimize}_{\theta}\;\left\{\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}y_{i}-\theta_{0}-\sum_{j=1}^{p}{(\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij})}\theta_{j}\Bigr{)}^{2}+\lambda\sum_{j=1}^{p}\theta_{j}\right\}\quad\mbox{ with$\theta_{j}\geq 0\;\forall j$}.$
    Select $\lambda$ by cross-validation.
  3. 3.
    The final model can be written as $\hat{\eta}(x)=\hat{\gamma}_{0}+\sum_{j=1}^{p}\hat{\gamma}_{j}x_{j}$, with $\hat{\gamma}_{j}=\hat{\beta}_{j}\hat{\theta}_{j}$, and $\hat{\gamma}_{0}=\hat{\theta}_{0}+\sum_{\ell=1}^{p}\hat{\beta}_{0\ell}\hat{\theta}_{\ell}$.
This procedure is computationally convenient: in Step 1 we can use efficient LOO formulas, and in Step 2 we can apply any efficient $\ell_{1}$ solver. Here we used the R language package glmnet. Specifically, we use the function cv.glmnet to estimate the lasso path parameter, and have all of the functionality of glmnet at our disposal. We provide a function cv.UniLasso in the R package uniLasso that implements this approach.
Note that we do not standardize the features before applying the non-negative lasso in Step 2; the univariate LOO fits are all on the scale of the response. From our knowledge of multiple linear regression, the first constraint — agreement between univariate and multivariate signs — may seem like an unreasonable restriction. However our belief is that in high-dimensions there are likely to be a multitude of different models that have about the same MSE as the “optimal” one chosen by the lasso. Hence it can make sense to choose one that is interpretable and sparser than that of lasso.
@@
These conjectures are borne out by our numerical studies. In simulations with varying problems sizes and SNR, and a number of real datasets (see Section 8), in almost every case uniLasso did no worse than the lasso in terms of out-of-sample MSE, and delivered a substantially sparser model. We note that Meinshausen (2012) studies sign-constrained least squares estimation for high-dimensional regression, which relates to our non-negativity constraint in our second step.
  1. Remark A.
@@
    The uniLasso procedure applies seamlessly to other GLMs as well the Cox model. Indeed, all of the models covered by the glmnet package are at our disposal. All we need are the separate fitted linear models and their LOO fit vectors in step 1, and then glmnet can be directly applied. The only challenge is find an (approximate) LOO formula for the GLM families. We give details of these computations in Section 16.
  2. Remark B.
@@
    UniLasso can be thought of as a version of stacked regression, or Stacking (Wolpert, 1992; Breiman, 1996). Stacking is a two-stage procedure for combining the predictions from a set of models (“learners”), in order to get improved predictions. It works as follows. A set of base models is trained on the training data; these models can be of the same type (e.g. gradient boosting) or different types (e.g., linear regression, decision trees, etc.). Each model generates predictions, capturing specific aspects of the relationship between the predictors and the target variable. A meta-model is then trained to combine the predictions of the base models into a final prediction. This meta-model learns how to weigh and integrate the LOO outputs of the base models to minimize the overall prediction error. UniLasso is a special case of stacking where the individual learners are simple univariate regressions.
  3. Remark C.
@@
    Why do we use the LOO univariate estimates $\hat{\eta}_{j}^{-i}=\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij}$ as features in step 2, instead of the univariate estimates $\hat{\eta}_{j}^{i}=\hat{\beta}_{0j}+\hat{\beta}_{j}x_{ij}$? [^1] In this case we could simply scale the features by their univariate slope coefficients and ignore the intercepts, which would be handled by the overall intercept in the model. In traditional stacking this is essential because the individual learners can be of very different complexity. In uniLasso, one would think that the learners (univariate regression) all have the same complexity so that the LOO versions are not needed. Indeed, the theory in Section 7 is based on the LOO estimates but holds equally well for the usual (non-LOO) estimates. However in practice we have found that the LOO estimates lead to greater sparsity and better performance, and hence we use them. With the use of the usual (non-LOO) univariate estimates in uniLasso, the resulting estimator is closely related to the adaptive lasso and we explore this connection in Section 3.
  4. Remark D.
@@
    The non-negative garotte (Breiman, 1995) is another closely related method. It minimizes $\sum_{i}(y_{i}-\sum_{j}c_{j}\hat{\beta}_{j}x_{ij})^{2}$ subject to $c_{j}\geq 0$ for all $j$, and $\sum c_{j}\leq s$, where $\hat{\beta}_{j}$ are the usual (multivariable) LS estimates. This is different from our proposal in an important way: our use of univariate LOO estimates in the first step. The univariate coefficients lead to materially different solutions and also allow application to the $p>n$ scenario.
@@ 2.2 UniLasso with no-regularization
We call this “uniReg” for univariate-guided regression and it represents an interesting alternative to the usual least squares estimates. For example, the non-negative constraint can still provide sparsity even though there is no $\ell_{1}$ penalty. If $p>n$ there can be multiple solutions; in this case we look for the sparsest solution by computing the limiting uniLasso solution as $\lambda\downarrow 0$. The standard error and distribution of the estimated coefficients can be estimated via the bootstrap. We study uniReg in detail in Section 9.
2.3 Two examples — one good, one bad
Homecourt. We generated data with 100 observations, 30 standard normal features with AR(1) correlation 0.8, 20% sparsity and non-negative coefficients in two stages, as follows:
$y^{\prime}$$\leftarrow$ $\sum_{j}x_{j}\beta_{j}+\sigma^{\prime}z^{\prime}$ (2)
From these data we compute the $p=30$ univariate least squares coefficients $\hat{\beta}_{j}^{uni}$ separately for each $j$, and then generate $y$ as
$y$$\leftarrow$ $\sum_{j}x_{j}\hat{\beta}_{j}^{uni}\beta_{j}+\sigma z.$ (3)
The variance terms $\sigma^{\prime}$ and $\sigma$ were chosen so that at both stages the SNR was 1. The idea here is that (3) roughly mimics the model fit by uniLasso in its second stage.
Figure 1: Results for homecourt example: CV and test set prediction error. The dashed vertical lines for each method correspond to the models chosen by CV.
@@ 2.3 Two examples — one good, one bad
Mean 1.098 1.077 7.660 4.790
se 0.005 0.005 0.309 0.142
TPR-lasso TPR-uniLasso FPR-lasso FPR-uniLasso
Mean 0.737 0.700 0.135 0.025
se 0.014 0.016 0.012 0.004
Table 2: Test MSE, support, TPR (True Positive Rate) and FPR (False Positive Rate) for 100 simulations from the setting of Figure 1.
Figure 1 shows the CV and test error curves for the lasso and uniLasso. We see that uniLasso has test error a little below that of lasso, with a smaller active set. Table 2 shows the result of 100 simulations from this setup. We see that the same pattern emerges, and uniLasso exhibits a slightly lower true positive rate and a much lower false positive rate than the lasso.
[…]
@@ 3 Relationship of uniLasso to the adaptive lasso
The adaptive lasso (Zou, 2006) is defined by
$\hat{\gamma}^{*}={\rm argmin}=\frac{1}{2}\sum_{i=1}^{n}(y_{i}-\gamma_{0}-\sum_{j=1}^{p}x_{ij}\gamma_{j})^{2}+\lambda\sum_{j=1}^{p}w_{j}|\gamma_{j}|$ (4)
where $w_{j}=1/|\hat{\gamma}_{j}|^{\nu}$. Here the vector $\hat{\gamma}$ is any root-$n$ consistent estimate of $\gamma$, for example the least squares estimates.
The original paper assumed $n>p$; with $p>n$, one cannot use the least-squares estimates. One possibility is to use an initial estimate (from say ridge or lasso), solve the adaptive lasso, and iterate. Another option is to use the univariate least-squares estimates; (Huang et al., 2008) consider a combination of these to achieve some theoretical guarantees.
Let $\{\hat{\beta}_{j}\}_{1}^{p}$ be the univariate least squares coefficients. Consider the adaptive lasso with weights $w_{j}=1/|\hat{\beta}_{j}|,\;j=1,\ldots,p$. Then it is easy to show that this procedure is equivalent to uniLasso with $\hat{\eta}_{j}^{i}=\hat{\beta}_{0j}+\hat{\beta}_{j}x_{ij}$ replacing their LOO versions $\hat{\eta}_{j}^{-i}=\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij}$ in Step 2, and removing the non-negativity constraint.
@@
$\min_{\theta}\sum_{i=1}^{n}(y_{i}-\theta_{0}-\sum_{j=1}^{p}\eta_{j}^{i}\theta_{j})^{2}+\lambda\sum_{j=1}^{p}|\theta_{j}|\quad\mbox{ s.t. $\theta_{j}\geq 0\;\forall j$}$ (5)
$\min_{\gamma}\sum_{i=1}^{n}(y_{i}-\gamma_{0}-\sum_{j=1}^{p}x_{ij}\gamma_{j})^{2}+\lambda\sum_{j=1}^{p}\frac{|\gamma_{j}|}{|\hat{\beta}_{j}|}\quad\mbox{ s.t.$\operatorname{sign}(\gamma_{j})=\operatorname{sign}(\hat{\beta}_{j})\;\forall j$}$ (6)
@@
The proof is in the Appendix. However our experiments with this version of adaptive lasso produced models that were not nearly as sparse as the ones using LOO. In fact, in Table 3 we see that enforcing sign constraints led to less sparse models than not enforcing sign constraints!